OpenStudy (anonymous):

10w+4w=-84 how is this one sloved @mitchal

OpenStudy (bibby):

combine the like terms (10w+4w)

OpenStudy (anonymous):

tht would be 14w

OpenStudy (bibby):

and so now we have 14*w=-84 solve for w

OpenStudy (anonymous):

14w=-84 and you are done... bibby..look the question again, were only combining like terms not solving the equation!

OpenStudy (bibby):

what are you talking about?

OpenStudy (anonymous):

youre solving the equation, were only supposed to combine like terms right babe?

OpenStudy (bibby):

hahahahahaha

OpenStudy (bibby):

that's right, babe

OpenStudy (anonymous):

i have to slove this one boo @mitchal

OpenStudy (anonymous):

ahhh sorry my aplogized

OpenStudy (anonymous):

its okay hehe

OpenStudy (anonymous):

i been doing to much combining like terms... I aplogized bibby

DivineSolar (divinesolar):

So simplify both sides of the equation.

DivineSolar (divinesolar):

which gives you 14w = -84

OpenStudy (anonymous):

so 14w=-84 and we have to divide 14 on both side

OpenStudy (anonymous):

so what -84 divide by 14?

OpenStudy (anonymous):

tht would be 6 right

DivineSolar (divinesolar):

Now you divide both sides of the equation by 14, \[\frac{ 14w }{ 14 } = \frac{ -84 }{ 14 }\]

DivineSolar (divinesolar):

yes

OpenStudy (anonymous):

negative 6 actually but youre close. ;)

DivineSolar (divinesolar):

Well -6

OpenStudy (anonymous):

so w=-6 is your answer ;)

DivineSolar (divinesolar):

That is obvious.

DivineSolar (divinesolar):

xD

OpenStudy (anonymous):

:p lol divine XD

OpenStudy (anonymous):

thank you @DivineSolar

OpenStudy (anonymous):

thank you boo @mitchal

OpenStudy (anonymous):

no problem baby XD @shay16

DivineSolar (divinesolar):

Sure.